OpenStudy (anonymous):

evaluate lim (x^2-1)/(sqrt(x)-1) x-> the answer in the book says 2

OpenStudy (anonymous):

lim x->1 x^2-1/ sqrt(x) -1 = lim x->1 ( sqrt(x) -1 )( sqrt(x) +1 ) / ( sqrt(x) -1 ) = lim x->1 ( sqrt(x) +1) pluggin in the values , we get = ( sqrt(1) +1) = 1 +1 =2 hope that helps !!! let me know if u r satisfied...

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

I dont understand the first step

OpenStudy (anonymous):

okay let me explain

OpenStudy (anonymous):

remember the formula a^2 - b^2 = (a - b) (a +b)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(sqrt(x)-1) (sqrt(x)+1) does that equal x-1?

OpenStudy (anonymous):

u got it ...cheers !!!!

OpenStudy (anonymous):

but the original problem has a numerator of x^2-1 not X-1

OpenStudy (anonymous):

the square root of x^2 is only x

OpenStudy (anonymous):

okay !!! then make it as (x -1)(x +1) which equals to (sqrt(x)-1) (sqrt(x)+1) (x +1) and now plug in the values

OpenStudy (anonymous):

but thats like we are getting 4

OpenStudy (anonymous):

I thought this one looked so simple

OpenStudy (anonymous):

r u sure that u read the statement correct?

OpenStudy (anonymous):

evaluate the indicated limit, if it exists. lol (x-1)/(sqrt(x)-1) I was wrong

OpenStudy (anonymous):

yes.......the limit does exist which is 4

OpenStudy (anonymous):

your original answer is right, Thank you very much for your help!

OpenStudy (anonymous):

4 or 2 ? :)

OpenStudy (anonymous):

2 because I had the problem wrong the numerator is (x-1)

OpenStudy (anonymous):

ahh !!!!!! finally we r there !!! gud luck with ur maths ..

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

no worries !!!!